OpenStudy (anonymous):

what is the eqation to the line parrallel to line 2x-y=-2 given point (6,0)?

OpenStudy (anonymous):

y = 2x + 2 y = 2x - 12

OpenStudy (jwt625):

2x-y=12

OpenStudy (jwt625):

there is only one possible answer.

OpenStudy (anonymous):

Yeah, the first eqn is just me rearranging the given equation into something useful. Second eqn is the answer. Same as yours.

OpenStudy (jwt625):

I see.

OpenStudy (anonymous):

whre does the -12 come from?

OpenStudy (anonymous):

You want the second equation to be parallel so the gradient needs to be maintained. The gradient being the 2x so you know the second eqn must look like y = 2x + b Where b still needs to be figured out. Sub in the given coordinate (6,0) into this new equation and solve for b. 0 = 2*(6) + b b = 0 - 12 b = -12

OpenStudy (anonymous):

Well to be padantic the gradient is just the 2 infront of the x.

OpenStudy (anonymous):

The general equation of a line is y = mx + b Where; m = gradient (or the slope of the line, rise/run, y/x) b = the y-intercept (the location on the graph where the line crossing the y axis) With x and y being input variables x and y. If a line is to run parallel to another line the slope has to be the same, otherwise they would cross at some point and hence not be parallel.